OpenStudy (anonymous):

Which of the following is true for the rational function shown in the graph below?

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (mathstudent55):

Look at the answers with asymptotes. Does either one make sense?

OpenStudy (anonymous):

nop i know it has to be either B or D

OpenStudy (mathstudent55):

Ok. Look at B. Can you factor the denominator?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Show me.

OpenStudy (mathstudent55):

That's the numerator.

OpenStudy (anonymous):

xD Whoops

OpenStudy (mathstudent55):

Look at the denominator of B. x^2 + 16 Can this denominator be zero? Is there any value of x that would make this denominator zero?

OpenStudy (mathstudent55):

Now quite: (x + 4)(x + 4) = (x + 4)^2 = x^2 + 8x + 16 You have x^2 + 16, without the middle term of 8x.

OpenStudy (mathstudent55):

(x - 4)(x - 4) doesn't work either: (x - 4)(x - 4) = x^2 - 8x + 16 Again, there is amiddle term here that your problem does not have.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

(x-2)(x-2)

OpenStudy (mathstudent55):

x^2 + 16 for x = 2: 2^2 + 16 = 4 + 16 = 20

OpenStudy (mathstudent55):

Ok. Let's look at B. The denominator cannot be factored. x^2 + 16 cannot be factored. Now think this way: The only way x^2 + 16 could equal zero is if x^2 = -16, since -16 + 16 = 0. But x^2 cannot equal a negative number. That means, there is no value of x that can make x^2 + 16 equal zero. Since x^2 + 16 is in the denominator, and the denominator cannot be zero, B has no asymptotes.

OpenStudy (mathstudent55):

Now look at D: The denominator is x^2 - 16. x^2 - 16 factors into (x + 4)(x - 4) That means when x = -4 or x = 4, x^2 - 16 equals zero. That means x = -4 and x = 4 are not allowed. That means there is an asymptote at x = -4 and an asymptote at x = 4.

OpenStudy (anonymous):

ok so d is the answer

OpenStudy (mathstudent55):

Right.